OpenStudy (briana.img):

Find the number of permutations of the first 9 letters of the alphabet taking 6 letters at a time.

OpenStudy (briana.img):

Is it like 9*8*7*6*5*4 ??? = 60,480

jimthompson5910 (jim_thompson5910):

You are correct. You can also use the formula \[\Large P(n,r) = \frac{n!}{(n-r)!}\]

OpenStudy (briana.img):

I was hoping to learn it lol

OpenStudy (briana.img):

aaaah okay! thank you!

jimthompson5910 (jim_thompson5910):

n! = n*(n-1)*...*3*2*1 example: 7! = 7*6*5*4*3*2*1